# Strong anomalous diffusion for free-ranging birds

### v1

@@ III Results
For the adults $s_{1}=1/\xi=0.63\pm 0.06$ resulting in $\xi=1.59\pm 0.17$, where the physical meaning of the exponent $\xi$ is related to the broad tail of the flight time distribution, see details in simulations below. Notably, the value $q_{s}=2.14\pm 0.18$ for the adults is close to the value of 2, anticipated for this general class of Lévy walk models [66]. However (4) does not account for the results of the fledglings, see Table 1, or for the concave $\lambda_{\ell}(q)$ at $t>5$ minutes. These findings motivate the bounded Lévy walk model developed in Sec. IV.1.
Importantly, the bi-linear dependence of the moment spectrum function $\lambda_{s}(q)$ entails that the PDF does not adhere to mono-scaling in a self-similar fashion, which would imply a linear dependence of $\lambda_{s}(q)$ with $q$ [37]. As the empirical data exhibit a bi-linear form of $\lambda_{s}(q)$, we expect two distinct scaling regimes, a rescaled (self-similar) PDF $t^{1/\xi}P(|\bm{x}|)$ versus $|\bm{x}|/t^{1/\xi}$ for smaller displacements (bulk) and a rescaled PDF $t^{\xi}P(|\bm{x}|)$ versus $|\bm{x}|/t$ for larger displacements (the tails of the PDF) [37]. Both rescaled regimes directly follow from (4), whereas the latter indicates a ballisitic scaling. Indeed, analysis of the PDF of the displacements of the Barn Owls shows good agreement with these predictions. The normalized PDFs for different times $t$ are shown in Figs. 2a and 2b. In panel 2a we plot PDFs of the normalized variable $|\bm{x}|/t^{1/\xi}$ for different times $t$, showing that the data at different times in the bulk of the PDF collapse onto universal curves. Here, $\xi=1.59$ is not directly fitted for the collapse, but rather obtained from analysis of the moments, see (4). At the tails of the distribution, the scaling $|\bm{x}|/t^{1/\xi}$ is no longer observed. In panel 2b we plot PDFs of the normalized variable $|\bm{x}|/t$, showing a collapse at the tails of the distribution. In the insets of both panels we plot a histogram of the upper $2\%$ of the displacement for each time $t$, highlighting that the ballistic-like scaling $|\bm{x}|/t$ leads to a collapse of the tail at larger displacements. This, together with the non-linearity of $\lambda_{s}(q)$, indicates that mono-scaling theories are not sufficient to describe the dynamics.
@@
Figure 2: Normalized PDFs for the displacement at different $t$, where different symbols correspond to different time windows – ranging between 32 and 600 seconds. Panel (a) displays the PDF of the normalized variable $|\bm{x}|/t^{1/\xi}$, while panel (b) shows the PDF of $|\bm{x}|/t$. In the insets, we plot a normalized histogram of the top $2\%$ of the distances with x-axis in linear scale. These two panels exhibit bi-scaling, the scaling in (a) for the bulk which corresponds to enhanced diffusion with $\xi=1.59$, see (4), and the scaling in (b) for the tail which displays ballistic transport.
@@ IV.1 Bounded Lévy walk
The first suggested model is bounded Lévy walk (BLW), serving as a “zero-order approach”, i.e., an effective model with a low-parameter phase space. The BLW is constructed to account for the finite home range of the animal and is based on the well-known Lévy walk model. Similarly to Lévy walk, the BLW is characterized by a power-law distribution of jump durations – the times it takes the walker to move to a new location – according to $\psi(\tau)=1/\tau^{1+\alpha}$ for $\tau>1$ and a constant velocity $v$ [8]. We define BLW by a constraint, disallowing traversal beyond a predefined radius $R$ relative to the initial position (see Appendix A), such that $R$ serves to define a finite home range. As it has been shown for unbounded Lévy walk that $\xi=\alpha$ [66, 37], we use $\xi=1.59\pm 0.17$ above (see below Eq. (4)) and set $\alpha=1.6$. We have also checked that direct flights of the adult Barn Owls follow power law distributed $\psi(\tau)$ (Fig. S3 [65]), and note that in the BLW model, due to the exclusion of steps beyond the predefined radius, the jump-time distribution is cutoff. In our simulations we set $v=8$ m/s, in agreement with empirical data (Appendix A). In Fig. 3a we plot an example of 4-hour long BLW for $\alpha=1.6$ and $R=\mathcal{R}_{0}=2.5$ km, where the latter is approximately the birds’ average maximum displacement, see Sec. III.
@@
Figure 3: Analysis of BLW simulations gives results similar to empirical data. (a) 4-hour simulated tracks, sampled every 4 seconds. (b) Log-log plot of the average absolute displacement moments $\langle|\bm{x}|^{q}\rangle/\mathcal{R}_{0}^{q}$ with $\mathcal{R}_{0}=2.5$ km representing the average maximum displacement in a single night. Data points for moment orders $q=1,1.5,2,2.5,3$ are shown, displaying two distinct scaling regimes for each moment $q$, with a transition at $\sim 5$ minutes. (c) $\lambda_{s}(q)$ as a function of $q$ for $t<5$ minutes (red triangles). The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively, where the slopes are denoted in the legends. (d) $\lambda_{\ell}(q)$ as a function of $q$ for $t>5$ minutes (red triangles). The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively.
@@
We next repeat the analysis of the empirical data for the BLW simulations for $\alpha=1.6$ and $R=2.5$ km, where the results are shown in Fig. 3 and summarized in Table 1. The BLW shares several key similarities with the empirical data (Fig. 1), including: (a) a shift in the moments at a characteristic timescale of $\sim$5 minutes (Fig. 3b); (b) convex moment spectrum function for $t<5$ minutes quantitatively similar to the empirical results for the adult Barn Owls (Fig. 3c, compare Fig. 1c); and (c) a transition between a convex and a concave moment spectrum function (Fig. 3d). Notably, for $t<5$ minutes, the BLW model closely aligns with adult Barn Owls in terms of $s_{1}$, $s_{2}$, and $q_{s}$. For $t>5$ minutes, $\ell_{1}$ is similar and both $\ell_{2}$ and $q_{\ell}$ are larger in the BLW model, compared to adults; however, the model still captures a concave $\lambda_{\ell}(q)$. When comparing fledglings to the BLW model, we find more discrepancies. For $t<5$ minutes, fledgling values for $s_{1}$, $s_{2}$, and $q_{s}$ are lower, indicating an overestimation by the BLW model. For $t>5$ minutes, we find a small difference in $\ell_{1}$ and a larger difference in $q_{\ell}$.
@@
In addition to results for specific values of $\alpha$ and $R$, we have checked that for $21<\alpha<$ 2, variations in $R$ – but not in $\alpha$ – modify the form of the long-time $\lambda_{\ell}(q)$. Similarly, variations in $R$ affect the typical timescale for transitioning from convex to concave (Appendix A). Thus, the results of the BLW model with $R=2.5$ km, suggests that the observed phenomena are a mark of a finite home range, whose scale is given here by the maximum displacement scale $\mathcal{R}_{0}=2.5$ km, see further discussion below. We have shown that this model can reproduce key features observed in the birds, providing a basic understanding of a possible underlying mechanism. Finally, we note that for the BLW, $q_{s}>2$ for all values of $R$ and $\alpha$ that we checked, in contrast to $q_{s}$ for the fledglings. This suggests that the BLW model does not as effectively represent fledgling movement compared to adults, although it still successfully models the transition from a convex $\lambda_{s}(q)$ to a concave $\lambda_{\ell}(q)$. In the following, we employ a more ecologically oriented model to better contextualize these phenomena in an ecological setting.
@@ IV.2 Multi-mode biased random-walk
We implement a multi-mode two-dimensional biased random-walk process for animal movement, similar to processes commonly used in the ecological literature [68, 69, 16, 27, 52, 70]. Our model consists of biased random walks embedded within discrete behavioral states with transition probabilities between states [53, 71, 19]. Each behavioral state within the model is characterized by a correlation time, $\tau_{i}$, and a mean velocity, $\nu_{i}$, where the subscript $i$ denotes different behavioral modes (e.g., feeding, resting, or evading predators). The transitions between these behavioral states are governed by a Markovian transition matrix ($M$), dictating the likelihood of an animal shifting from one behavior to another. Further details of the simulations are given in Appendix B, and in what follows we limit ourselves to two modes: commutes and local area-restricted searches. An example of a 4-hour multi-mode simulation, with two distinct modes is given in Fig. 4a. We note that here a bounding radius is not directly incorporated, although the transitions between modes has been reported to act as an inhibitor of large displacements [70].
Although the model is parameter-rich, estimation of the parameters can be guided by empirical data. Here we estimate the parameters by segmentation of the data into two modes: searches and commutes (see Methods and Ref. [27]). Following the segmentation, we can derive a correlation time for each mode by examining the autocorrelation function of each mode [72]; the mean velocity is directly calculable from the displacements between adjacent relocation’s; and the transition matrix $M$ is inferred from the observed transitions between modes. This approach yields a direct estimation of the parameters, providing a basis to explore the model’s behavior in relation to the empirical findings, see Table 2 and Appendix B for further details.
@@
Table 2: Simulation Parameters for multi-mode OU simulation based on segmentation of the datasets. $\tau_{i}$ are the correlation times, $v_{i}$ are the velocities, and $M$ is the probability matrix to transition between states at any given time step.
Figure 4: Analysis of Multi-mode simulations shows results similar to empirical data. (a) 4-hour simulated tracks. (b) Log-log plot of the average absolute displacement moments $\langle|\bm{x}|^{q}\rangle/\mathcal{R}_{0}^{q}$ with $\mathcal{R}_{0}=2.5$ km representing the average maximum displacement of Barn Owls in a single night. Data points for moment orders $q=1,1.5,2,2.5,3$ are shown, displaying two distinct scaling regimes for each moment $q$, with a transition at $\sim 5$ minutes. (c) $\lambda(q)$ as a function of $q$. The red triangles represent the values over short time intervals of the simulations. The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively, see legends. (f) $\lambda(q)$ as a function of $q$ for long time intervals ($>5$ minutes, red triangles). The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively.
@@
V Discussion

### v2

@@ III Results
For the adults $s_{1}=1/\xi=0.63\pm 0.06$ resulting in $\xi=1.59\pm 0.17$, where the physical meaning of the exponent $\xi$ is related to the broad tail of the flight time distribution, see details in Sec. IV below. However (4) does not account for the results of the fledglings, see Table 1, or for the concave $\lambda_{\ell}(q)$ at $t>5$ minutes. These findings motivate the bounded Lévy walk model developed in Sec. IV.1.
As discussed in Sec. I, the bi-linear dependence of the moment spectrum function $\lambda_{s}(q)$ entails that the PDF does not adhere to mono-scaling in a self-similar fashion, which would imply a linear dependence of $\lambda_{s}(q)$ with $q$ [37]. As the empirical data exhibit a bi-linear form of $\lambda_{s}(q)$, we expect two distinct scaling regimes, a rescaled (self-similar) PDF $t^{1/\xi}P(|\bm{x}|)$ versus $|\bm{x}|/t^{1/\xi}$ for smaller displacements (bulk) and a rescaled PDF $t^{\xi}P(|\bm{x}|)$ versus $|\bm{x}|/t$ for larger displacements (the tails of the PDF) [37]. Both rescaled regimes directly follow from (4), whereas the latter indicates a ballisitic scaling. As such, this bi-scaling has important statistical consequences for the asymptotic behavior of the underlying probability densities.
Analysis of the PDF of the displacements of the adult Barn Owls shows good agreement with these predictions. The normalized PDFs for different times $t$ are shown in Figs. 2a and 2b. In panel 2a we plot PDFs of the normalized variable $|\bm{x}|/t^{1/\xi}$ for different times $t$, showing that the data at different times in the bulk of the PDF collapse onto universal curves. Here, $\xi=1.59$ is not directly fitted for the collapse, but rather obtained from analysis of the moments, see (4). At the tails of the distribution, the scaling $|\bm{x}|/t^{1/\xi}$ is no longer observed. In panel 2b we plot PDFs of the normalized variable $|\bm{x}|/t$, showing a collapse at the tails of the distribution. In the insets of both panels we plot a histogram of the upper $2\%$ of the displacement for each time $t$, highlighting that only the ballistic-like scaling $|\bm{x}|/t$ in Fig. 2b, leads to a collapse of the tail at large displacements. Thus, we find that bi-scaling with ballistic motion for the tails of the propagator and $|\bm{x}|/t^{1/\xi}$ for the central part, are applicable to describe the dynamics.
@@
Figure 2: Normalized PDFs for the displacement at different $t$, where different symbols correspond to different time windows – ranging between 32 and 600 seconds (lower blue circles to upper green pluses, respectively). Panel (a) displays the PDF of the normalized variable $|\bm{x}|/t^{1/\xi}$, while panel (b) shows the PDF of the ballistic scaling $|\bm{x}|/t$. In the insets, we plot a normalized histogram of the top $2\%$ of the distances with x-axis in linear scale, where the axis labels of the main panels apply also to the insets. These two panels exhibit bi-scaling, the scaling in (a) for the bulk which corresponds to enhanced diffusion with $\xi=1.59$, see Eq. (4), and the scaling in (b) for the tail which displays ballistic scaling.
@@ IV.1 Bounded Lévy walk
The first suggested model is bounded Lévy walk (BLW), serving as a “zero-order approach”, i.e., an effective model with a low-parameter phase space. The BLW is constructed to account for the finite home range of the animal and is based on the well-known Lévy walk model. Similarly to Lévy walk, the BLW is characterized by a power-law distribution of jump durations – the times it takes the walker to move to a new location – according to $\psi(\tau)=1/\tau^{1+\alpha}$ for $\tau>1$ and a constant velocity $v$ [8]. We define BLW by constraining movement within a fixed radius $R$ from the initial position; any sampled jump exceeding this boundary is resampled until a valid one is obtained (see Appendix A), such that $R$ serves to define a finite home range. As it has been shown for unbounded Lévy walk that $\xi=\alpha$ [67, 37], we use $\xi=1.59\pm 0.17$ above (see below Eq. (4)) and set $\alpha=1.6$. We have also checked that direct flights of the adult Barn Owls follow power law distributed $\psi(\tau)$ (Fig. S2 [66]), and note that in the BLW model, due to the exclusion of steps beyond the predefined radius, the jump-time distribution is cutoff. In our simulations we set $v=8$ m/s, in agreement with empirical data (Appendix A). In Fig. 3a we plot an example of 4-hour long BLW for $\alpha=1.6$ and $R=\mathcal{R}_{0}=2.5$ km, where the latter is approximately the birds’ average maximum displacement, see Sec. III. We stress that the BLW model has two parameters, $\alpha$ and $R$, which are chosen to match $\xi=1.59\pm 0.17$ and $\mathcal{R}_{0}=2.5$ km, respectively.
@@
Figure 3: Analysis of BLW simulations with parameters $\alpha=1.6$ and $R=2.5$ gives results similar to empirical data. (a) 4-hour simulated tracks, sampled every 4 seconds. (b) Log-log plot of the average absolute displacement moments $\langle|\bm{x}|^{q}\rangle/\mathcal{R}_{0}^{q}$ with $\mathcal{R}_{0}=2.5$ km representing the average maximum displacement in a single night. Data points for moment orders $q=1,1.5,2,2.5,3$ are shown, displaying two distinct scaling regimes for each moment $q$, with a transition at $\sim 5$ minutes. (c) $\lambda_{s}(q)$ as a function of $q$ for $t<5$ minutes (red triangles). The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively, where the slopes are denoted in the legends. (d) $\lambda_{\ell}(q)$ as a function of $q$ for $t>5$ minutes (red triangles). The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively.
@@
We next repeat the analysis of the empirical data for the BLW simulations for $\alpha=1.6$ and $R=2.5$ km, and the results are shown in Fig. 3 and summarized in Table 1. The BLW shares several key similarities with the empirical data (Fig. 1), including: (a) a shift in the moments at a characteristic timescale of $\sim$5 minutes (Fig. 3b); (b) convex moment spectrum function for $t<5$ minutes quantitatively similar to the empirical results for the adult Barn Owls (Fig. 3c, compare Fig. 1c); and (c) a transition between a convex and a concave moment spectrum function (Fig. 3d). Notably, for $t<5$ minutes, the BLW model closely aligns with adult Barn Owls in terms of $s_{1}$, $s_{2}$, and $q_{s}$. For $t>5$ minutes, $\ell_{1}$ is similar and both $\ell_{2}$ and $q_{\ell}$ are larger in the BLW model, compared to adults; however, the model still captures a concave $\lambda_{\ell}(q)$. When comparing fledglings to the BLW model, we find more discrepancies. For $t<5$ minutes, fledgling values for $s_{1}$, $s_{2}$, and $q_{s}$ are lower, indicating an overestimation by the BLW model. For $t>5$ minutes, we find a small difference in $\ell_{1}$ and a larger difference in $q_{\ell}$.
@@
In addition to results for specific values of $\alpha$ and $R$, we have checked that for $1<\alpha<2$, variations in $R$ – but not in $\alpha$ – modify the form of the long-time $\lambda_{\ell}(q)$. Similarly, variations in $R$ affect the typical timescale for transitioning from convex to concave (Appendix A). Thus, the results of the BLW model with $R=\mathcal{R}_{0}=2.5$ km, suggests that the observed phenomena are a mark of a finite home range, whose scale is given by the maximum displacement scale $\mathcal{R}_{0}=2.5$ km, see further discussion below. Finally, we note that for the BLW, $q_{s}>2$ for all values of $R$ and $\alpha$ that we checked, in contrast to $q_{s}$ for the fledglings. This suggests that the BLW model does not as effectively represent fledgling movement compared to adults, although it still successfully models the transition from a convex $\lambda_{s}(q)$ to a concave $\lambda_{\ell}(q)$. In the following, we employ a more ecologically oriented model to better contextualize these phenomena in an ecological setting.
@@ IV.2 Multi-mode biased random-walk
We implement a multi-mode two-dimensional biased random-walk process for animal movement, similar to processes commonly used in the ecological literature [69, 70, 16, 27, 53, 71]. Our model consists of a random walker that transitions between discrete behavioral states, where each state corresponds to a different mode of stochastic evolution [54, 72, 19]. Each behavioral state within the model is fully characterized by a correlation time, $\tau_{i}$, and a mean velocity, $\nu_{i}$, where the subscript $i$ denotes different behavioral modes (e.g., feeding, resting, or evading predators). The transitions between these behavioral states are governed by a Markovian transition matrix ($M$), dictating the likelihood of an animal shifting from one behavior to another. Further details of the simulations are given in Appendix B, and in what follows we limit ourselves to two modes: commutes and local area-restricted searches. An example of a 4-hour multi-mode simulation, with two distinct modes is given in Fig. 4a. We note that here a bounding radius is not directly incorporated, although the transitions between modes has been reported to act as an inhibitor of large displacements [71].
Although the model is parameter-rich, estimation of the parameters can be guided by empirical data. Here we estimate the parameters by segmentation of the data into two modes: searches and commutes (see Sec. II.2 and Ref. [27]). Following the segmentation, we can derive a correlation time for each mode by examining the autocorrelation function of that mode [73]; the mean velocity is directly calculable from the displacements between adjacent relocation’s; and the transition matrix $M$ is inferred from the observed transitions between modes. This approach yields a direct estimation of the parameters, providing a basis to explore the model’s behavior in relation to the empirical findings. The resulting parameters are summarized in Table 2, and see Appendix B for more details.
@@
Table 2: Simulation Parameters for multi-mode OU simulation based on segmentation of the datasets. $\tau_{i}$ are the correlation times, $v_{i}$ are the velocities, and $M$ is the probability matrix to transition between states.
Figure 4: Analysis of Multi-mode simulations shows results similar to empirical data. (a) 4-hour simulated tracks. (b) Log-log plot of the average absolute displacement moments $\langle|\bm{x}|^{q}\rangle/\mathcal{R}_{0}^{q}$ with $\mathcal{R}_{0}=2.5$ km representing the average maximum displacement of Barn Owls in a single night. Data points for moment orders $q=1,1.5,2,2.5,3$ are shown, displaying two distinct scaling regimes for each moment $q$, with a transition at $\sim 5$ minutes. (c) $\lambda_{s}(q)$ as a function of $q$. The red triangles represent the values over short time intervals of the simulations. The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively, see legends. (f) $\lambda_{\ell}(q)$ as a function of $q$ for long time intervals (>5>5 minutes, red triangles). The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively.
@@
V Ecological Significance
As shown above, observing SAND in adults and fledglings uncovered age-specific scaling regimes. At the same time, our analysis also highlighted commonalities: bi-linear scaling of the moment spectrum function and a characteristic temporal scale of approximately 5 minutes for both adults and fledglings. This shared timescale illustrates that while the structure and flexibility of movement modes differ ontogenetically, both age groups are constrained by a common temporal scale.
To understand the significance of this 5-minute timescale, we note that the commuting velocity for these birds is between $8$ and $9$ m/s (Table 2) and the maximum displacement per night is $\sim 2.5$ km as detailed above. Factoring both the maximum displacement and the velocity, a bird flying at 898-9 m/s will cover 2.5 km at approximately 5 minutes. Based on this observation and the models detailed above we offer two interpretation of this timescale. The first is a coupling between the individual and the environment as supported by the BLW model: due to the finite home range of the bird, straight long-range flights are limited to $\sim 5$ minutes, above which interactions with the edges of the home range become important and inhibit the increase of the displacement. This explanation is supported by direct analysis of the home range of the birds, where in Appendix C and Fig. S4 [66] we show that indeed for the adult Barn Owls, the home range during breeding is 232-3 km wide. On the other hand, we have found that – on average – the fledglings display significantly larger home ranges than the adults (Fig. S4 [66]).
Another closely related explanation for the 5-minute timescale is based on the multi-mode correlated-velocity model, and does not require incorporating explicit boundaries. In the multi-mode model, mechanical or preferential behavior of the bird limits its commuting behavior, resulting in SAND at short times and successfully producing a transition from convex $\lambda_{s}(q)$ to concave $\lambda_{\ell}(q)$ without an explicit home range. Notably, the multi-mode model can result in a relatively short (average) daily maximum displacement but would allow for larger excursions, as found for the fledglings. Here, an important finding is that the straightforward parameterization of the model – without detailed fine-tuning – accurately reflects our qualitative results as well as the quantitative differences between adults and fledglings, suggesting that our two-state multi-mode model captures the dynamics to a significant degree.
V.1 Differences between age groups
The differences in movement parameters between fledglings and adults – namely, shorter correlation times during localized search ($\tau_{1}$), longer correlation times during commutes ($\tau_{2}$), and lower average velocities in fledglings compared to adults ($\nu_{1}$ and $\nu_{2}$) – are consistent with documented patterns of age-specific foraging proficiency in birds. Juveniles of several species exhibit lower capture efficiency, less selective habitat use, and reduced foraging success due to inexperience, underdeveloped motor coordination, and limited spatial memory [56, 74, 75, 76, 77]. The more sharply delineated behavioral modes observed in fledglings (stronger separation in $\tau_{1}$, $\tau_{2}$, $v_{1}$, and $v_{2}$) may reflect a compensatory strategy: discrete, simplified movement regimes may help manage energetic costs and reduce cognitive demands while foraging proficiency is still developing. This interpretation is supported by studies noting that juveniles often engage in more stereotyped search patterns, rely more heavily on trial-and-error or local enhancement, and are more vulnerable to suboptimal foraging decisions (e.g., overexploitation of poor patches or inefficient prey handling) [56, 75, 77]. It is also consistent with the larger home ranges observed for fledglings in our data and in other species, where adults tend to maintain narrower home ranges, while juveniles continue to explore and sample broader areas for varying reasons [78, 79, 76, 77].
From an adaptive perspective, the distinct convex $\lambda_{s}(q)$ and lower crossover point $q_{s}$ observed in fledglings could reflect limited flexibility in switching between exploratory and exploitative behaviors, possibly due to a need for extended learning or increased sensitivity to predation risk. Adults, by contrast, exhibit smoother transitions between movement modes, higher velocities, and more integrated search-commute patterns, likely facilitated by experience or by spatial knowledge of resource distributions. These differences may therefore reflect a maturation process that enhances both spatial and behavioral efficiency, ultimately conferring a fitness advantage through improved prey detection, reduced energy expenditure, and increased reproductive provisioning success [56]. Our analysis and simulations provide quantitative support for these patterns; however, we note that the biological interpretations offered here are based primarily on parallels with findings from other species. Our results thus provide a biologically grounded framework for generating testable hypotheses about how developmental constraints shape foraging dynamics.
V.2 Stochastic modeling
Further relating the two stochastic models to the data, we suggest the following explanation for the emergence of SAND: in a heterogeneous landscape, an animal may alternate between an extensive commuting mode – search for resource-rich patches – and an intensive, area-concentrated searching mode – search for prey within a local patch [55]. Indeed, a major effort in the literature is to distinguish between local area-restricted searches and commutes [80, 81]. Thus, an analysis of unsegmented trajectories results in a mixture of directed flights and long stationary periods, which might resemble BLW dynamics [8, 55]. As discussed above, this mixture occurs at short times scales, up to $\sim 5$ minutes, either due to interactions with the environment – at long time-scales commutes are limited by the animal’s home-range – or due to properties of the behavioral states such as energy-consumption or food-availability. The BLW model thus acts as a simple model for SAND, simplifying the more ecologically inclined multi-mode dynamics. We further stress that while the BLW does not aim to replicate specific behavioral processes, it captures the key statistical features of strong anomalous diffusion and serves as a useful reference point for interpreting the more complex, biologically grounded, OU simulations. In this sense, our findings offer another example of how Lévy walk-like movement patterns can emerge from underlying behavioral structure [82, 83, 84], helping to explain the widespread reports of Lévy walks in ecological systems [85].
Indeed, Lévy walks have been widely discussed in the ecological literature [23, 8, 86, 83], in the human mobility literature [87, 88, 89, 84], and in the context of biological movement, including Dictyostelium discoideum amoeba [90] and run-and-tumble bacteria [91], where movement patterns have been approximated by alternating sequences of relatively straight trajectories and reorientations [8]. While these models are highly simplified representations of complex behavior, they are used to construct workable descriptions of motion in diverse systems [13]. They have also contributed to theoretical discussions about the scales of movement, behavioral stationarity, and how to interpret trajectory data across spatial and temporal resolutions [55, 11]. As shown above, one characteristic of Lévy walks – and specifically our BLW model – is the emergence of strong anomalous diffusion (SAND), suggesting that SAND may be a broadly relevant phenomenon.
Finally, our stochastic models highlight an important feature of movement analysis: the sensitivity of displacement statistics to temporal scale. The observed crossover around five minutes in $\lambda_{t}(q)$ reflects ecological timescales intrinsic to the system and is consistent across our BLW and multi-mode OU process. Crucially, the agreement between simulations and data arises from parameterizations that align these timescales, such as the bounding radius $R$ in the BLW model and the transition matrix $\mathbf{M}$ in the OU model. Modifying any of the model parameters may shift the crossover point between temporal regimes. Although a full exploration of this sensitivity is beyond the scope of the present study, future work should examine how high moments in multi-mode models, including the OU process and other run-and-tumble dynamics [92, 93, 94], respond to variation in temporal and spatial scales.
VI Discussion